# On the interpretation of XRISM X-ray measurements of turbulence in the intracluster medium: a comparison with cosmological simulations

### v1

@@ 4 Conclusions
The physics of turbulence in the weakly collisional intracluster medium is expected to be more complicated than that in fluid numerical simulations, where an idealised collisional plasma is assumed. In this article we have investigated whether the recently observed properties of turbulent gas motions inferred via X-ray spectroscopy by XRISM Collaboration et al. (2025a) are alrady in tension with the idealised view from numerical simulations on large sales () and have already started to pinpoint aspect driven by the collisionless nature of the intracluster medium, which are expected to arise on smaller scales (e.g. Kunz et al., 2011; Brunetti & Lazarian, 2011; Santos-Lima et al., 2014; Brunetti & Jones, 2014; Rincon et al., 2016; St-Onge & Kunz, 2018; St-Onge et al., 2020; Kempf & Rincon, 2025).
In particular, using a typical high-resolution cosmological simulation of a Coma-like cluster of galaxies (Sec.2), we have shown that:
@@
spectroscopic X-ray estimates of the ICM turbulence along the LOS are likely to be biased towards the X-ray brightest X-ray portion of the ICM volume, which produce an underestimate (overall by a factor on average, compared to the energy weighted turbulent dispersion along the LOS) due to two factors: the limited range of scales of the turbulent spectrum sampled in this way, and the limited radial extent from the cluster centre which can be probed by X-ray observations (which are biased towards the innermost dense cluster regions).
@@
The fact that turbulent energy and X-ray emission only loosely correlate in the ICM, explains why the total turbulent kinetic pressure support in our Coma-like cluster of galaxies is of the total pressure (or up to if the full unfiltered kinetic energy is considered), even if our simulated X-ray line broadening is consistent with the observation by XRISM Collaboration et al. (2025a). The actual turbulent pressure support in the ICM can plausibly higher than what is estimated from X-ray spectroscopy, although it is not possible to derive a simple correction factor, due to the large measured scatter in the relation between the true and the X-ray weighted turbulent dispersion along the LOS (see also, e.g. Roncarelli et al., 2018; Biffi et al., 2022, for valuable numerical work on this issue).
The velocity structure function of gas motions in the simulated ICM is well in line with Kolomogorov turbulence in a stratified atmosphere, and consistent with the last two decades of numerical simulation. Nevertheless, our tests show that if the VSF is estimated just using a limited set of FOV from X-ray observations, the resulting VSF is too affected by statistical uncertainties to be a reliable proxy of turbulent properties.
@@
In summary, we suggest that the fluid turbulence typically produced by cosmological simulations of clusters in the literature is compatible with the latest XRISM measurements in Coma (and possibly in other clusters), once that the realistic degree of non-homogeneity of the ICM, the multi-scale nature of the ICM turbulent flows and the biases inherent to X-ray spectroscopy are taken into account. This appear enough to release the tension between the XRISM data and current numerical simulations, and it shows that a turbulent spectrum much steeper than Kolmogorov is not required by current data.
@@ acknowledgements
FV has been partially supported by Fondazione Cariplo and Fondazione CDP, through grant n° Rif: 2022-2088 CUP J33C22004310003 for the ”BREAKTHRU” project. F.V. acknowledges the CINECA award ”IsCc4_FINRADG” under the ISCRA initiative, for the availability of high-performance computing resources and support.

### v2

@@ 4 Conclusions
The physics of turbulence in the weakly collisional ICM is expected to be more complicated than that in fluid numerical simulations, where an idealised collisional plasma is assumed. In this article we have investigated whether the recently observed properties of turbulent gas motions inferred via X-ray spectroscopy by Coma_XRISM are alrady in tension with the idealised view from numerical simulations on large sales () and have already started to pinpoint aspects driven by the collisionless nature of the ICM, which are expected to arise on smaller scales (e.g. 2011MNRAS.410.2446K; bl11; 2014ApJ...781...84S; bj14; 2016PNAS..113.3950R; 2018ApJ...863L..25S; 2020JPlPh..86e9003S; 2025A&A...694A..25K).
In particular, using a homogeneous and cluster-volume filling model of turbulence, the VSF and X-ray line width measured by XRISM require a turbulent spectrum that is much steeper than a Kolmogorov, possibly pin-pointing either a very large effective viscosity in the ICM or a situation where large scale motions generated by recent mergers have not yet cascaded down to small scales. Using a high-resolution cosmological simulation of a Coma-like cluster (Sec.2), we have shown that :
the spatially patchy, not homogeneous and multi-scale turbulence in the simulation generates a VSF and X-ray line-widths integrated along the LOS that are both consistent with the XRISM observations, thus releasing the request of a turbulent spectrum that is much steeper than a Kolmogorov. This also shows that the interpretation of XRISM data is very sensitive to the assumed turbulent model.
@@
X-ray lines integrated along the LOS are sensitive to the X-ray brightest regions. This biases the effect of line broadening due to turbulence to an extent that depends on the properties of the turbulence itself in the integrated volume. In our simulation we find that the simulated widths of XRISM X-ray lines are biased low by with respect to the turbulent velocity dispersion on scale that is measured along the LOS. This bias arises from two factors : the limited range of scales of the turbulent spectrum sampled in this way, and the limited radial extent from the cluster centre which can be probed by X-ray observations (which are biased towards the innermost dense cluster regions).
@@
The pressure support in our simulated cluster from turbulence on filtered scales of 300-600 kpc is estimated in the range , which is times larger than that estimated from the widths of X-ray lines obtained from the simulation, which are consistent with those measured by Coma_XRISM. The actual turbulent pressure support in the ICM can plausibly higher than what is estimated from X-ray spectroscopy, although it is not possible to derive a simple correction factor, due to the large measured scatter in the relation between the true and the X-ray weighted turbulent dispersion along the LOS (see also, e.g. 2018A&A...618A..39R; 2022A&A...663A..76B; 2025arXiv250614441L, for valuable numerical work on this issue).
@@
In summary, we suggest that the fluid turbulence typically produced by cosmological simulations of galaxy clusters (including our ones and the majority of those in the literature) is compatible with the latest XRISM measurements in Coma, and possibly in other clusters, provided that the realistic degree of non-homogeneity of the ICM, the multi-scale nature of the ICM turbulent flows and the biases inherent to X-ray spectroscopy are fully taken into account.
More in general, our results highlights the importance of advanced numerical simulations to support the interpretation of the data from X-ray calorimeters in the future.
@@ acknowledgements
FV has been partially supported by Fondazione Cariplo and Fondazione CDP, through grant n° Rif: 2022-2088 CUP J33C22004310003 for the ”BREAKTHRU” project. GB acknowledges partial support from INAF Theory Grant ”Theory and simulations of non-thermal phenomena in galaxy clusters and beyond”. F.V. acknowledges the CINECA award ”IscrB_CREW” under the ISCRA initiative, for the availability of high-performance computing resources and support. We wish to thank our reviewer for the valuable comments on our draft, and S. Ettori, D. Eckert, T. Lebeau, A. Ignesti, F. Rincon, M. Markevitch, V. Biffi, J. Zuhone and I. Zhuravleva for helpful comments and discussons. The authors wish to declare that of their original manuscript has been produced using AI of any sort.
Appendix A Comparison with radial profiles of the Coma cluster
We compared the radial profiles of gas density, and gas entropy (defined as where is the gas temperature) from our Coma-like simulated cluster, with the available observational constraints for the real Coma. This is shown by the two panels in Fig.6. The simulated density profile compares reasonably well with -model parametrisation recently obtained by churazov21 using X-ray observations by the eROSITA satellite, and also with the pointed deep X-ray exposures of two narrow selections along the E+NE and NW+W directions in the Coma cluster, take by simionescu13 using SUZAKU. The lower panel shows that the average gas entropy profile of our simulated cluster is also fairly in line with the most recent constrains obtained by simionescu13 using SUZAKU. The reasonably good comparison with these two constraints for the real Coma cluster is reassuring as it shows that the gas matter distribution (and its related X-ray emission) from our model is realistic, and also that the entropy stratification of our simulated cluster (which in turns is related to buoyancy and gas mixing as a function of radius) is in line with observations.
Appendix B Comparison with idealised turbulence
In order to quantify the importance of the turbulent model, here we present the comparison between the line width obtained by integrating the simulation along the LOS and that obtained using a homogeneous and volume filling Kolmogorov turbulent model, similar to the procedure adopted in Coma_XRISM and in 2025arXiv251021918E. To calculate the homogenous model, we initialised a spherically symmetric -model gas density profiles with parameters of the Coma cluster (e.g. churazov21, and references therein) for a volume using cells, and we generated a random 3-dimensional velocity field in Fourier domain, assuming it follows the Kolmogorov scaling () from the largest to the smallest available grid scales. In order to obtain a random Gaussian distribution of velocities, we randomly extracted vectors components from the Rayleigh distribution. The 3-dimensional velocity field within the volume generated in such a way is renormalised to exactly match the kinetic energy budget, within an equivalent volume centred on the cluster centre, in our cosmological simulation. Figure 7 shows the distribution of gas density and velocity along the line of sight for two middle planes in the two datasets, while Fig. 8 shows the comparison of the 3-dimensional velocity power spectra of the two datasets. The latter confirms the trend of the VSF of Fig.2: the spectrum of gas motions in the simulated cluster approximately follows the Kolmogorov spectrum up to , while it steepens for scales (see Sec.3 for a discussion).
In Figure 9 we show the distribution of the X-ray weighted gas velocity dispersion along random lines of sight through the two simulated datasets. The distribution derived from the cosmological simulation of our cluster is clearly peaked at a smaller value, , while the distribution derived from the idealised turbulent model peaks at the higher value of .
This simple experiment (although limited to the central fraction of the simulated volume, and not to the full depth of the lines of sight used in Fig.4) very well illustrates that, for a realistic dynamical model of turbulence in the ICM, a smaller typical broadening of X-ray lines is to be expected, compared to a more idealised model of turbulence, despite the total kinetic energy of the two models is the same. The key factor here is the realistic level of intermittency of turbulence naturally produced by the multi-scale accretions onto cluster simulated in cosmology: turbulence has in this case a smaller filling factor compared to fully homogenous models which assume a large injection scale and stationary driving. For the same total kinetic energy budget, a more intermittent turbulence distribution produces on average a smaller velocity dispersion along the line of sight. Conversely, if two models produce on average the same broadening of X-ray lines along the line of sight, homogenous model is biased to be characterised by significantly smaller kinetic energy budget. Finally, with an additional test (green line) we show the distribution of X-ray weighted gas velocity dispersions obtained in a ”mixed” model, in which we use the -model density distribution to compute the X-ray emission, and combine this with the 3-dimensional velocity field of the cosmological simulation. The distribution obtained in this way is quite close the one obtained using the density distribution from the simulation, and it shows that the key factor here is the intermittency in the turbulent velocity field, and not in the density field.